OpenStudy (anonymous):

The millennium falcon with a mass of 850 kg with a velocity of 101 m/s. What is the momentum of the ship? I got 85,850 kgm/s a. The velocity of the Millennium Falcon is increased to 159.0 m/s in a time of 2.7 seconds. What is the momentum now after the change in velocity? b. What is the Impulse for the Falcon?

OpenStudy (anonymous):

I dont get a.) or b.) :(

OpenStudy (anonymous):

momentum = mass * velocity M1=850Kg*101m/s =85850Kgm/s now velocity is increased to 159m/s so new momentum(M2)= 850Kg*159m/s=135150Kgm/s now impulse=change of momentum/time =(M2-M1)/t=49300/2.7=18259.259N :)

OpenStudy (anonymous):

there's only 1 mass in this equation :O

OpenStudy (anonymous):

so what that is not masses that are momentums:)

OpenStudy (anonymous):

soo i do (135,150 - 85850) which are the momentums

OpenStudy (anonymous):

what's the Impulse? o.O

OpenStudy (anonymous):

The two momentums are 85,850 for the first one and 135,150 for the second one

OpenStudy (anonymous):

wait nvm lols